2. Franklin Construction entered into a fixed-price contract to build a freeway-connecting ramp for $58 million. Construction costs incurred in the first year were $48 million and estimated remaining costs to complete at the end of the year were $29 million.

a. How much gross profit or loss will Franklin recognize in the first year if it recognizes revenue over time according to percentage of completion method? (Enter your answer in millions.)

b.
How much gross profit or loss will Franklin recognize in the first year applying the completed contract method? (Enter your answer in millions.)
Answer #1

2. a. Total estimated cost = 48+29 = $77 million

Loss = Contract price - Total estimated cost

= 58 - 77 = $19 million

Net revenue recognized = (contract price x % of completion) - Loss

= (58*48/77) - 19 = $17.15 million

Gross loss recognize in the first year = Construction cost incurred - Net revenue recognized

= 48 - 17.15 = $30.85 million

b. Gross loss to be recognized in the first year under the completed contract method = Contract price - Estimated cost

= $58 - 77

= $19 million